# Semantic Equitable Clustering: A Simple, Fast and Effective Strategy for Vision Transformer

### v1

@@ Abstract
The Vision Transformer (ViT) has gained prominence for its superior relational modeling prowess. However, its global attention mechanism’s quadratic complexity poses substantial computational burdens. A common remedy spatially groups tokens for self-attention, reducing computational requirements. Nonetheless, this strategy neglects semantic information in tokens, possibly scattering semantically-linked tokens across distinct groups, thus compromising the efficacy of self-attention intended for modeling inter-token dependencies. Motivated by these insights, we introduce a fast and balanced clustering method, named Semantic Equitable Clustering (SEC). SEC clusters tokens based on their global semantic relevance in an efficient, straightforward manner. In contrast to traditional clustering methods requiring multiple iterations, our method achieves token clustering in a single pass. Additionally, SEC regulates the number of tokens per cluster, ensuring a balanced distribution for effective parallel processing on current computational platforms without necessitating further optimization. Capitalizing on SEC, we propose a versatile vision backbone, SecViT. Comprehensive experiments in image classification, object detection, instance segmentation, and semantic segmentation validate to the effectiveness of SecViT. Remarkably, SecViT attains an impressive 84.2% image classification accuracy with only 27M parameters and 4.4G FLOPs, without the need for for additional supervision or data. Code will be available at https://github.com/qhfan/SecViT.
@@ 1 Introduction
Figure 2: Top-1 accuracy v.s. FLOPs on ImageNet-1K of resent SOTA models. SecViT outperforms all the counterparts in all settings.
[…]
@@ 2 Related works
The Vision Transformer (ViT) [13] is considered a powerful visual architecture due to its excellent visual modeling capabilities. Many works have improved the Vision Transformer, including enhancing its training efficiency and reducing its computational cost [36, 12, 47, 26, 62]. DeiT uses distillation loss and incorporates extensive data augmentation methods into the ViT training process. This increases the model’s data utilization efficiency, allowing it to train an effective ViT with less data. Hierarchical structures represented by PVT [51, 52, 45, 16, 54] reduce the number of tokens in global attention by downsampling the keys and values (KV), thereby keeping the computational cost of global attention within an acceptable range. In addition to using downsampling to reduce the number of tokens, some methods directly prune tokens based on their importance, retaining important tokens while discarding or merging unimportant ones [15, 42]. This reduces the number of tokens and subsequently lowers the computational cost of the model. Another highly representative approach is to group all tokens such that each token can only attend to tokens within its own group [36, 12, 62, 11, 35]. This method also significantly reduces the computational cost of self-attention.
@@
Most grouping-based attention mechanisms perform grouping based on spatial structure [36, 12, 48, 11, 35]. Specifically, the Swin-Transformer [36] divides all tokens into equally sized windows based on their spatial positions, where each token can only attend to tokens within its own window. This significantly reduces the model’s computational cost. In contrast, CSWin [12] divides tokens into a crisscrossing structure, which also reduces the model’s computational cost. In addition to dividing tokens into small windows along the spatial dimension, DaViT [11] also splits channels into multiple groups along the channel dimension. Unlike the above methods that only consider positional information for grouping, DGT [35] takes semantic information into account by using k-means clustering to group the queries.
@@ 3.1 Overall Architecture
The overall architecture of SecViT is shown in Fig. 3(a). For each input image $x\in\mathbb{R}^{3\times H\times W}$, we first process it by a patch embedding to downsample it. Then we obtain the tokens with a shape of $C_{1}\times\frac{H}{4}\times\frac{W}{4}$. Similar to previous work [36, 51, 12], we adopt a hierarchical structure to process tokens. Our model consists of four stages with downsampling factors of $\frac{1}{4}$, $\frac{1}{8}$, $\frac{1}{16}$, and $\frac{1}{32}$, respectively. This structural design facilitates downstream tasks, such as object detection, in constructing feature pyramids.
A SecViT block is composed of three modules: Conditional Positional Encoding (CPE) [7], Semantic Equitable Clustering based Self-Attention (SECA), and classical Feed-Forward Network (FFN) [49]. A complete SecViT block can be represented by Eq. 1:
$\centering\begin{split}&X={\rm CPE(}X_{in}{\rm)}+X_{in},\\ &Y={\rm SECA(}{\rm LN(}X{\rm))}+X,\\ &Z={\rm FFN(}{\rm LN(}Y{\rm))}+Y.\end{split}\@add@centering$ (1)
For each block, the input tensor $X_{in}\in\mathbb{R}^{C\times H\times W}$ is fed into the CPE to introduce the positional information. Then, The Self-Attention based on the Semantic Equitable Clustering (SEC) is employed to serve as the token mixer. The final FFN is utilized to integrate channel-wise information of tokens.
@@ 3.2 Semantic Equitable Clustering
Figure 3: (a) The overall architecture of our Semantic Equitable Clustering Vision Transformer. (b) The composition of each SecViT block. (c) Illustration of our Semantic Equitable Clustering. SEC determines the clustering rules for tokens by calculating the cosine similarity between all tokens and the global token.
@@
To address these two issues, we first discard the use of multiple cluster centers and instead calculate the distance between each token and a single center. Based on each token’s distance to this center, we divide the tokens into different intervals. Then, to ensure that our chosen center contains the most comprehensive semantic information, we directly use the result of average pooling of all tokens as the center token. This is because, in most vision foundation models, the output of the average pool is assumed to contain the richest semantic information and is thus used for classification [36, 12, 7, 14].
ToMe [2] indicates that the $K$ contains the richest information from each token. Therefore, we choose the $K$ as the basis for clustering. Specifically, the process for determining the cluster center is shown in Eq. 2:
$\centering\begin{split}Q=W_{Q}X,&K=W_{K}X,V=W_{V}X,\\ &k_{c}={\rm Avg}(K).\end{split}\@add@centering$ (2)
@@
The obtained $Q^{*}$, $K^{*}$, and $V^{*}$ from the previous step have been sorted based on their distances to the cluster center. At this point, we directly group them, so tokens with similar distances to the cluster center are classified into the same cluster. This allows us to directly control an equal number of tokens in each cluster. This process can be clearly illustrated in Fig. 3(c) and denoted as follows:
$\centering\begin{split}Q_{c}=Q_{c\times N_{c}:\left(c+1\right)N_{c}}^{*},K_{c}=K_{c\times N_{c}:\left(c+1\right)N_{c}}^{*},K_{c}=K_{c\times N_{c}:\left(c+1\right)N_{c}}^{*}.\end{split}\@add@centering$ (4)
@@
where $N_{c}$ is the basic token number of each cluster for equal partition and $c$ is the index of the cluster
@@
Based on the above steps, we have completed the clustering process that captures semantic information in the image with minimal sorting cost. Moreover, compared to K-means, we have achieved equi-partitioning of each cluster. After clustering is completed, we apply standard Self-Attention to the tokens within each cluster (SECA), thereby completing the interaction of information between tokens:
$\centering Y_{c}=\text{Attn}\left(Q_{c},K_{c},V_{c}\right).\@add@centering$ (5)
After the completion of SECA, to further enhance the model’s ability to capture local information, we employ a local context enhancement module to model local information, where LCE [62, 12] is a simple depth-wise convolution:
$\centering\begin{split}X={\rm SECA}(X)+{\rm LCE}(W_{V}X),\end{split}\@add@centering$ (6)
@@ 4 Experiments
We conducted experiments on a wide range of vision tasks, including image classification on ImageNet-1k [10], object detection and instance segmentation on COCO [33], and semantic segmentation on ADE20K [61]. All models can be trained with 8 A100 80G GPUs.
4.1 ImageNet Classification
Settings.
We train our models from scratch on ImageNet-1k [10]. For a fair comparison, we follow the training strategy in DeiT [47], with classification loss serving as the only supervision. The maximum rates for increasing stochastic depth [23] are set to 0.1, 0.15, 0.4, and 0.5 for SecViT-T, SecViT-S, SecViT-B, and SecViT-L, respectively.
@@ 4.1 ImageNet Classification
SecViT-T 15 2.5 82.6
small model $\sim 4.5$G Swin-T [36] 29 4.5 81.3
CrossViT-15 [4] 27 5.8 81.5
RVT-S [38] 23 4.7 81.9
@@
SecViT-S 27 4.4 84.2
Cost Model Parmas (M) FLOPs (G) Top1-acc (%)
base model $\sim 9.0$G Swin-S [36] 50 8.7 83.0
ConvNeXt-S [37] 50 8.7 83.1
CrossFormer-B [53] 52 9.2 83.4
@@
SE-CoTNetD-152 [31] 56 26.5 84.6
SecViT-B 55 9.7 85.0
large model $\sim 18.0$G DeiT-B [47] 86 17.5 81.8
Swin-B [36] 88 15.4 83.3
LITv2 [40] 87 13.2 83.6
CrossFormer-L [53] 92 16.1 84.0
Ortho-L [25] 88 15.4 84.2
CSwin-B [12] 78 15.0 84.2
@@
SecViT-L 96 18.1 85.6
@@
Comparison with SOTA.
We compare our SecViT with numerous state-of-the-art models, the results are shown in Tab.1. SecViT consistently outperforms preceding models across all scales. Notably, SecViT-S attains a Top1-accuracy of 84.2% with a mere 15M parameters and 2.5G FLOPs. For larger models, SecViT-L achieves a Top1-accuracy of 85.6% with 96M parameters and 18.1G FLOPs.
Model Parmas (M) FLOPs (G) Throughput (imgs/s) Top1-acc (%)
Swin-T [36] 29 4.5 1723 81.3
SEC-Swin-T 29 4.9 1482 83.6(+2.3)
Swin-S [36] 50 8.8 1062 83.0
SEC-Swin-S 50 9.3 804 84.8(+1.8)
Model Parmas (M) FLOPs (G) Throughput (imgs/s) Top1-acc (%)
FasterViT-0 [19] 31 3.3 3551 82.1
SEC-FasterViT-0 25 3.6 3218 82.6(+0.5)
FasterViT-1 [19] 53 5.3 2619 83.2
SEC-FasterViT-1 41 5.6 2486 83.8(+0.6)
Table 2: Strict comparison with the baslines on ImageNet-1K classfication. The speed of the models are measured on A100 GPU with the batch size of 64.
Strict Comparison with baselines.
We select two baselines: a general-purpose backbone, Swin-Transformer [36], and a efficiency-oriented backbone, FasterViT [19] to make a comparison with our SecViT. In the comparison models (SEC-Swin and SEC-FasterViT), we merely substitute the attention mechanism in the original Swin-Transformer and FasterViT with our Semantic Equitable Clustering based Self-Attention and without introducing any other modules (such as CPE, Conv Stem, etc.). As shown in Tab. 2, the simple replacement of the attention mechanism yields significant advantages in both performance and efficiency. Specifically, SEC-Swin achieves or even surpasses a 2.0% improvement over Swin in Tiny and Small model sizes. Meanwhile, SEC-FasterViT attains higher accuracy than FasterViT while utilizing fewer parameters.
4.2 Object Detection and Instance Segmentation
Settings.
We utilize MMDetection [6] to implement Mask-RCNN [21], Cascade Mask R-CNN [3], and RetinaNet [32] to evaluate the performance of the SecViT. For Mask R-CNN and Cascade Mask R-CNN, we follow the commonly used “3×+3$\times$+3 × + MS” setting, and for Mask R-CNN and RetinaNet, we apply the “1×1$\times$1 ×” setting. Following previous works [36, 12, 62], during training, we resize the images such that the shorter side is 800 pixels while keeping the longer side within 1333 pixels.
Backbone Params (M) FLOPs (G) Mask R-CNN 3×3$\times$3 ×+MS
$AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$
Focal-T [57] 49 291 47.2 69.4 51.9 42.7 66.5 45.9
NAT-T [18] 48 258 47.8 69.0 52.6 42.6 66.0 45.9
@@ 4.2 Object Detection and Instance Segmentation
CSWin-S [12] 54 342 50.0 71.3 54.7 44.5 68.4 47.7
SecViT-B 73 368 52.5 73.4 57.6 46.1 70.6 49.8
Backbone Params (M) FLOPs (G) Cascade Mask R-CNN 3×3$\times$3 ×+MS
$AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$
@@
Ortho-S [25] 81 755 52.3 71.3 56.8 45.3 68.6 49.2
HorNet-T [43] 80 728 52.4 71.6 56.8 45.6 69.1 49.6
CSWin-T [12] 80 757 52.5 71.5 57.1 45.3 68.8 48.9
SecViT-S 83 738 53.9 72.6 58.6 46.7 70.1 51.1
Swin-S [36] 107 838 51.9 70.7 56.3 45.0 68.2 48.8
@@
SecViT-B 111 847 55.1 73.9 59.8 47.6 71.5 51.6
Table 3: Comparison with other backbones using "$3\times+\mathrm{MS}$“ schedule on COCO.
Results.
Tab. 3 and Tab. 4 show the results of SecViT with different detection frameworks. The results show that SecViT performs better than its counterparts in all comparisons. For “3×+3$\times$+3 × +MS” schedule, SecViT outperforms the recent InternImage [50] for +2.2 box AP and +1.7 mask AP with the framework of Mask R-CNN. Regarding the Cascade Mask R-CNN, SecViT still has very good performance. As for “1×1$\times$1 ×” schedule, SecViT also demonstrate very impressive performance.
SecViT-T 34 221 47.7 69.5 52.6 42.9 66.4 46.1 24 202 45.6 66.8 48.9 28.8 49.8 60.7
Swin-T [36] 48 267 43.7 66.6 47.7 39.8 63.3 42.7 38 248 41.7 63.1 44.3 27.0 45.3 54.7
CMT-S [16] 45 249 44.6 66.8 48.9 40.7 63.9 43.4 44 231 44.3 65.5 47.5 27.1 48.3 59.1
STViT-S [24] 44 252 47.6 70.0 52.3 43.1 66.8 46.5
FAT-B3 [14] 49 47.6 69.7 52.3 43.1 66.4 46.2 39 45.9 66.9 49.5 29.3 50.1 60.9
BiFormer-S [62] 47.8 69.8 52.3 43.2 66.8 46.5 45.9 66.9 49.4 30.2 49.6 61.7
SecViT-S 45 259 49.7 70.8 54.7 44.2 68.1 48.0 35 240 48.1 69.3 51.8 31.3 53.0 63.6
Swin-S [36] 69 359 45.7 67.9 50.4 41.1 64.9 44.2 60 339 44.5 66.1 47.4 29.8 48.5 59.1
InternImage-S [50] 69 340 47.8 69.8 52.8 43.3 67.1 46.7
CSWin-S [12] 54 342 47.9 70.1 52.6 43.2 67.1 46.2
STViT-B [24] 70 359 49.7 71.7 54.7 44.8 68.9 48.7
SecViT-B 73 368 51.2 72.7 56.4 45.4 69.7 49.0 63 349 48.9 70.2 52.7 32.0 53.6 64.6
Swin-B [36] 107 496 46.9 69.2 51.6 42.3 66.0 45.5 98 477 45.0 66.4 48.3 28.4 49.1 60.6
MPViT-B [28] 95 503 48.2 70.0 52.9 43.5 67.1 46.8 85 482 47.0 68.4 50.8 29.4 51.3 61.5
CSwin-B [12] 97 526 48.7 70.4 53.9 43.9 67.8 47.3
InternImage-B [50] 115 501 48.8 70.9 54.0 44.0 67.8 47.4
SecViT-L 115 546 51.9 73.4 57.0 46.1 70.4 49.6 105 527 49.9 71.2 53.7 33.0 54.5 66.1
Table 4: Comparison to other backbones using "1×1$\times$1 ד schedule on COCO.
4.3 Semantic Segmentation
Settings.
We utilize Semantic FPN [27] and UperNet [55] to validate our SecViT’s performance, implementing these frameworks via MMSegmentation [8]. We follow the PVT’s [51] training settings for Semantic FPN and train the model for 80k iterations. All models use an input resolution of $512\times 512$, and during testing, the image’s shorter side is resized to 512 pixels. Different from FPN, the UperNet is trained for 160K iterations and follows Swin’s [36] settings. We employ the AdamW optimizer with a weight decay of 0.01 and a 1500 iteration warm-up.
Semantic FPN
Backbone Params(M) FLOPs(G) mIoU(%)
PVTv2-B1 [52] 18 34 42.5
FAT-B2 [14] 17 32 45.4
EdgeViT-S [39] 17 32 45.9
SecViT-T 18 34 47.2
DAT-T [54] 32 198 42.6
CSWin-T [12] 26 202 48.2
Shuted-S [44] 26 183 48.2
FAT-B3 [14] 33 179 48.9
SecViT-S 29 177 49.5
DAT-S [54] 53 320 46.1
RegionViT-B+ [5] 77 459 47.5
UniFormer-B [29] 54 350 47.7
CSWin-S [12] 39 271 49.2
SecViT-B 57 288 50.7
DAT-B [54] 92 481 47.0
CrossFormer-L [53] 95 497 48.7
CSWin-B [12] 81 464 49.9
SecViT-L 99 471 52.0
UperNet
Backbone Params(M) FLOPs(G) mIoU(%)
DAT-T [54] 60 957 45.5
NAT-T [18] 58 934 47.1
InternImage-T [50] 59 944 47.9
MPViT-S [28] 52 943 48.3

### v2

# Semantic Equitable Clustering: A Simple and Effective Strategy for Clustering Vision Tokens

@@ Abstract
The Vision Transformer (ViT) has gained prominence for its superior relational modeling prowess. However, its global attention mechanism’s quadratic complexity poses substantial computational burdens. A common remedy spatially groups tokens for self-attention, reducing computational requirements. Nonetheless, this strategy neglects semantic information in tokens, possibly scattering semantically-linked tokens across distinct groups, thus compromising the efficacy of self-attention intended for modeling inter-token dependencies. Motivated by these insights, we introduce a fast and balanced clustering method, named Semantic Equitable Clustering (SEC). SEC clusters tokens based on their global semantic relevance in an efficient, straightforward manner. In contrast to traditional clustering methods requiring multiple iterations, our method achieves token clustering in a single pass. Additionally, SEC regulates the number of tokens per cluster, ensuring a balanced distribution for effective parallel processing on current computational platforms without necessitating further optimization. Capitalizing on SEC, we propose a versatile vision backbone, SECViT. Comprehensive experiments in image classification, object detection, instance segmentation, and semantic segmentation validate the effectiveness of SECViT. Moreover, SEC can be conveniently and swiftly applied to multimodal large language models (MLLM), such as LLaVA, to serve as a vision language connector, effectively accelerating the model’s efficiency while maintaining unchanged or better performance.
@@ 1 Introduction
Figure 2: Left: Top-1 accuracy v.s. FLOPs on ImageNet-1K of recent SOTA models. Right: Comparison among different vision language connectors on LLaVA-1.5
Building upon Semantic Equitable Clustering (SEC), we introduce the Semantic Equitable Clustering Vision Transformer (SECViT), a versatile vision backbone that is adaptable to a wide spectrum of downstream tasks. As demonstrated in Fig. 2, SECViT exhibits significant performance improvements compared to previous state-of-the-art (SOTA) models. Impressively, SECViT attains an accuracy of 84.3% utilizing merely 4.6GFLOPS, without the need for additional training data or supervision. This superior performance is maintained across different model scales. Furthermore, SECViT proves its proficiency in downstream tasks, including but not limited to, object detection, instance segmentation, and semantic segmentation.
Beyond vision tasks, we also apply SEC to multimodal large language models (MLLM) such as LLaVA-1.5 [39] to serve as an efficient vision language connector. Specifically, we use SEC to cluster the vision tokens, and then merge all the tokens at corresponding positions within each cluster into a single token. Experiments demonstrate that this approach significantly enhances the efficiency of LLaVA-1.5 while improving the model’s performance.
@@ 2 Related Works
The Vision Transformer (ViT) [13] is considered a powerful visual architecture. Many works have improved the Vision Transformer, including enhancing its training efficiency and reducing its computational cost [41, 12, 46, 27, 66]. DeiT [46] uses distillation loss and incorporates extensive data augmentation methods into the ViT training process. Hierarchical structures represented by PVT [48, 49, 44, 18, 54] reduce the number of tokens in global attention by downsampling the keys and values (KV), thereby low the computational cost. In addition to them, some methods directly prune tokens based on their importance, retaining important tokens [43, 35]. This reduces the number of tokens and subsequently lowers the computational cost of the model. Another highly representative approach is to group all tokens such that each token can only attend to tokens within its own group [41, 12, 66, 11, 40]. This method also significantly reduces the computational cost of self-attention.
@@
Most grouping-based attention mechanisms perform grouping based on spatial structure [41, 12, 47, 11, 40]. Specifically, the Swin-Transformer [41] divides all tokens into equally sized windows based on their spatial positions, where each token can only attend to tokens within its own window. This significantly reduces the model’s computational cost. In addition to dividing tokens into small windows along the spatial dimension, DaViT [11] also splits channels into multiple groups along the channel dimension. Unlike the above methods that only consider positional information for grouping, DGT [40] takes semantic information into account by using k-means clustering to group the queries.
Vision Language Connector.
The vision language connector is a critical component in MLLMs [39, 3, 26]. It aligns vision tokens with language tokens. Typical vision language connectors include MLP [39], Resampler [1], C-Abstractor [3], and others. Although MLP performs well, it introduces a significant number of vision tokens, which hampers the model’s efficiency. On the other hand, connectors like Resampler improve the model’s efficiency, but at the cost of reduced performance. Unlike these methods, our proposed SEC consider the semantic information of each token and significantly enhances the model’s efficiency while maintaining its performance.
Figure 3: (a) Illustration of SECViT (b) Applying SEC to vision language connector. (c) Illustration of Semantic Equitable Clustering for ViT and vision language connector.
@@ 3.1 Overall Architecture
The overall architecture of SECViT is shown in Fig. 3(a). SECViT consists of four stages with downsampling factors of $\frac{1}{4}$, $\frac{1}{8}$, $\frac{1}{16}$, and $\frac{1}{32}$, respectively. This structural design facilitates downstream tasks, such as object detection, in constructing feature pyramids. A SECViT block is composed of three modules. For each block, the input tensor $X_{in}\in\mathbb{R}^{C\times H\times W}$ is fed into the CPE to introduce the positional information. Then, The Self-Attention based on the Semantic Equitable Clustering (SEC) is employed to serve as the token mixer. The final FFN is utilized to integrate channel-wise information of tokens.
Beyond the design of the backbone, we also utilize SEC in the design of the vision language connector in MLLM [39]. For the vision tokens output by ViT, we use SEC to cluster the tokens. For each position corresponding to a cluster, we use attentive pooling to merge them into a single token, thereby reducing the number of vision tokens. The process is shown in Fig. 3(b).
@@ 3.2 Semantic Equitable Clustering
To address these two issues, we first discard the use of multiple cluster centers and instead calculate the distance between each token and a single center. Based on each token’s distance to this center, we divide the tokens into different intervals. Then, to ensure that our chosen center contains the most comprehensive semantic information, we directly use the result of average pooling of all tokens as the center token. This is because, in most vision foundation models, the output of the average pool is assumed to contain the richest semantic information and is thus used for classification [41, 12, 6, 14]. Specifically, the process for determining the cluster center is shown in Eq. 1:
$\centering\begin{split}Q=W_{Q}X,&K=W_{K}X,V=W_{V}X,\\ &k_{c}={\rm Pool}(K).\end{split}\@add@centering$ (1)
@@
The obtained $Q^{*}$, $K^{*}$, and $V^{*}$ from the previous step have been sorted based on their distances to the cluster center. For the design of vision backbone, we directly group them, so tokens with similar distances to the cluster center are classified into the same cluster. This allows us to directly control an equal number of tokens in each cluster. This process can be clearly illustrated in Fig. 3(c) and denoted as follows:
$\centering\begin{split}Q_{m}=Q^{*}[{m\times N:\left(m+1\right)N}],\\ K_{m}=K^{*}[{m\times N:\left(m+1\right)N}],\\ V_{m}=V^{*}[{m\times N:\left(m+1\right)N}].\end{split}\@add@centering$ (3)
@@
where $N$ is the basic token number of each cluster for equal partition and $m$ is the index of the cluster
@@
Based on the above steps, we have completed the clustering process that captures semantic information in the image with minimal sorting cost. Moreover, compared to k-means, we have achieved equi-partitioning of each cluster. After clustering is completed, we apply standard Self-Attention to the tokens within each cluster, thereby completing the interaction of information between tokens:
$\centering Y_{m}=\text{Attn}\left(Q_{m},K_{m},V_{m}\right).\@add@centering$ (4)
For the design of vision language connector, we group the tokens according to their similarity, and the tokens within each group are interleaved, as shown in Eq. 5:
$\centering\begin{split}Q_{n}=Q^{*}[n:N:L],\\ K_{n}=K^{*}[n:N:L],\\ V_{n}=V^{*}[n:N:L].\end{split}\@add@centering$ (5)
in which $L$ is the token’s sequence length, $n$ is the index of group tokens. $N$ is the basic token number of each cluster. After obtaining the token groups, we perform pooling on $Q$ to effectively reduce the number of tokens input to the LLM, with each group’s output becoming a single token, as shown in Eq 6.
$Y_{n}={\rm Attn}({\rm Pool}(Q_{n}),K_{n},V_{n}).$ (6)
3.3 Difference between SEC and EViT.
We use the most representative example, EViT [35], to illustrate the differences between SEC and other methods based on the similarity between the global token and other tokens.
Pruning v.s. Clustering. Most similarity-based methods, such as EViT, are pruning methods, where tokens with low similarity to the [cls] token are merged during the forward process, thereby reducing the number of tokens and decreasing computational cost. In contrast, our proposed SECViT employs a clustering-based approach, performing attention operations within each cluster.
The role of the [cls] token. In methods like EViT, the [cls] token serves as a measure of importance of a token. Each token computes its similarity to the [cls] token, with higher similarity tokens deemed more important. The less important tokens are abandoned. In contrast, in SEC, the [cls] token (obtained by average pooling over all tokens) measures similarity between tokens. Each token computes its similarity score to the [cls] token; tokens with similar scores are considered to be more similar and grouped into one cluster. Attention is calculated only within the same cluster.
Models Params (M) FLOPs (G) Throughput (imgs/s) Acc $AP^{b}$ $AP^{m}$ mIoU
DeiT-S 22 4.6 3204 79.8 44.5 40.1 43.0
EViT-DeiT-S (keeprate=0.9) 22 4.0 3428 79.8 not suit not suit not suit
SEC-DeiT-S (num_cluster=2) 22 4.3 3226 80.6 (+0.8) 47.9 (+3.4) 43.1 (+3.0) 48.0 (+5.0)
SEC-DeiT-S (num_cluster=4) 22 4.1 3412 80.5 (+0.7) 47.7 (+3.2) 42.7 (+2.6) 47.5 (+4.5)
SEC-DeiT-S (num_cluster=8) 22 3.9 3528 80.1 (+0.3) 46.7 (+2.2) 42.0 (+1.9) 46.4 (+3.4)
Table 1: Comparison of EViT [35] and SEC. Inference speed are measured on the A100 GPU.
Different adaptability to vision models/MLLMs. In pruning methods like EViT, during the model’s forward pass, the number of tokens gradually decreases. Although reducing tokens does not impact classification tasks, it prevents the feature map from being restored to its original shape. This makes it difficult for EViT to be directly used with classic frameworks like SemanticFPN for downstream dense prediction tasks . In SEC, we simply group the tokens without changing their quantity, thereby preserving the integrity of the feature map. This ensures that SEC can be easily applied to downstream tasks. In Tab. 1, we use DeiT [46] as the baseline. Without introducing any other tricks, just by applying SEC to DeiT, we form SEC-DeiT and compare it with EViT-DeiT. We conducted image classification, object detection, instance segmentation and semantic segmentation based on SEC-DeiT-S. SEC not only accelerates the model but also enhances its performance. Beyond the vision tasks, we also compare the EViT with our SEC on MLLM, details can be found in the Tab. 8.
@@ 4 Experiments
We first make strict comparison with hierarchical/plain baselines. Then we conduct experiments on a wide range of vision tasks for SECViT, including image classification, object detection, instance segmentation, and semantic segmentation. We also verify the role of SEC in MLLM based on LLaVA-1.5 [39]. More details, experiments, and comparison of models’ efficiency can be found in the Appendix.
4.1 SEC for vision models
Model Params (M) FLOPs (G) Throughput (imgs/s) Acc $AP^{b}$ $AP^{m}$ mIoU
DeiT-S [46] 22 4.6 3204 79.8 44.5 40.1 43.0
EViT-DeiT-S (keeprate=0.9) 22 4.0 3428 79.8 not suit not suit not suit
SEC-DeiT-S (num_cluster=4) 22 4.1 3412 80.5 (+0.7) 47.7 (+3.2) 42.7 (+2.6) 47.5 (+4.5)
DeiT-B 86 17.6 1502 81.8
SEC-DeiT-B 86 14.8 1682 82.4 (+0.6)
Swin-T 29 4.5 1723 81.3 43.7 39.8 44.5
SEC-Swin-T 29 4.8 1482 83.8 (+2.5) 48.3 (+4.6) 43.4 (+3.6) 49.3 (+4.8)
Swin-S 50 8.8 1006 83.0 45.7 41.1 47.6
SEC-Swin-S 50 9.2 804 85.0 (+2.0) 50.2 (+4.5) 44.7 (+3.6) 51.3 (+3.7)
Table 2: Comparison with Hierarchy/Plain baselines. Inference speed are measured on the A100 GPU.
Model Params (M) FLOPs (G) Method Pretrain epoch Acc(%)
Swin-B [41] 88 15.4 Supervised 83.5
ConvNeXt V2-B [53] 88 15.4 Supervised 84.3
SEC-Swin-B 88 16.2 Supervised 85.3
Swin-B [41] 88 15.4 SimMIM [56] 800 84.0(+0.5)
ConvNeXt V2-B [53] 88 15.4 FCMAE [53] 800 84.6(+0.3)
SEC-Swin-B 88 16.2 SimMIM [56] 800 85.9(+0.6)
Table 3: Comparison with baselines on self-supervised setting.
@@ 4.1 SEC for vision models
RMT-T [15] 14 2.5 82.4
SECViT-T 15 2.5 82.7
small model $\sim 4.5$G PS-ViT-B14 [61] 21 5.4 81.7
DVT-T2T-ViT-19 [52] 39 6.2 81.9
@@
SECViT-S 27 4.6 84.3
base model $\sim 9.0$G ConvNeXt-S [42] 50 8.7 83.1
@@
FAT-B4 [14] 52 9.3 84.8
SECViT-B 57 9.8 85.2
large model $\sim 18.0$G CrossFormer-L [50] 92 16.1 84.0
@@
RMT-L [15] 95 18.2 85.5
SECViT-L 101 18.2 85.7
XL model $\sim 35.0$G ConvNeXt-L [42] 198 34.4 84.3
CoAtNet-3 [10] 168 34.7 84.5
MaxViT-L [47] 212 43.9 85.1
GC ViT-L [22] 201 32.6 85.7
SECViT-XL 205 36.4 86.3
@@
RMT-T [15] 33 218 47.1 68.8 51.7 42.6 65.8 45.9 23 199 45.1 66.2 48.1 28.8 48.9 61.1
SECViT-T 34 221 47.8 69.5 52.5 43.0 66.7 46.3 24 202 45.8 66.8 49.2 29.1 49.8 60.9
MLLA-T [20] 44 255 46.8 69.5 51.5 42.1 66.4 45.0
STViT-S [25] 44 252 47.6 70.0 52.3 43.1 66.8 46.5
RMT-S [15] 46 262 49.0 70.8 53.9 43.9 67.8 47.4 36 244 47.8 69.1 51.8 32.1 51.8 63.5
SECViT-S 45 262 49.9 70.9 54.7 44.6 68.3 47.7 35 240 48.4 69.4 52.0 31.3 53.3 63.8
InternImage-S [51] 69 340 47.8 69.8 52.8 43.3 67.1 46.7
MLLA-S [20] 63 319 49.2 71.5 53.9 44.2 68.5 47.2
STViT-B [25] 70 359 49.7 71.7 54.7 44.8 68.9 48.7
SECViT-B 76 371 51.5 72.9 56.7 45.4 69.9 48.7 63 349 49.3 70.3 52.9 32.0 53.8 64.8
CSwin-B [12] 97 526 48.7 70.4 53.9 43.9 67.8 47.3
InternImage-B [51] 115 501 48.8 70.9 54.0 44.0 67.8 47.4
MLLA-B [20] 115 502 50.5 72.0 55.4 45.0 69.3 48.6
SECViT-L 119 550 52.0 73.5 57.3 46.3 70.6 49.8 105 527 50.2 71.4 53.9 33.2 54.5 66.3
Table 5: Comparison to other backbones using ”1×1$\times$1 ד schedule on COCO.
Backbone Params (M) FLOPs (G) $AP^{b}$ $AP^{b}_{50}$ $AP^{b}_{75}$ $AP^{m}$ $AP^{m}_{50}$ $AP^{m}_{75}$
Mask R-CNN 3×3$\times$3 ×+MS
@@
MLLA-S [20] 63 319 50.5 71.8 55.2 44.9 69.1 48.2
RMT-B [15] 73 373 52.2 72.9 57.0 46.1 70.4 49.9
SECViT-B 75 371 52.8 73.6 57.7 46.4 70.8 49.9
Cascade Mask R-CNN 3×3$\times$3 ×+MS
@@
RMT-S [15] 83 741 53.2 72.0 57.8 46.1 69.8 49.8
SECViT-S 83 741 54.1 72.8 58.6 47.0 70.3 51.0
@@
RMT-B [15] 111 852 54.5 72.8 59.0 47.2 70.5 51.4
SECViT-B 114 849 55.4 74.1 59.9 47.8 71.7 51.7
Table 6: Comparison with other backbones using ”$3\times+\mathrm{MS}$“ schedule on COCO.
Model Semantic FPN 80K Upernet 160K
Params (M) FLOPs (G) mIoU (%) Params (M) FLOPs (G) mIoUss (%)
PVTv2-B1 [49] 18 136 42.5
VAN-B1 [19] 18 140 42.9
RMT-T [15] 17 136 46.4
SECViT-T 18 136 47.2 44 894 48.8
StructViT-S [28] 26 271 46.9
MogaNet-S [34] 29 189 47.7 55 946 49.2